A roller coaster which weighs 640 lb starts from rest at point A and begins to coast down the track. If the frictional force is 5.0lb, how fast will the coaster be going at point B, which is 100 ft down the track and 30 ft below point A, and at point C, which is 50 ft farther along the track and 20 ft higher than point B?


1
Expert's answer
2022-03-02T10:07:04-0500

"v_B=v_A+\sqrt{2gh_{C}}-\sqrt{\frac{2Fl_B}m}=17~\frac{ft}s,"

"v_C=\sqrt{v_B^2+2gh_C}-\sqrt{\frac{2Fl_C}m}=19~\frac{ft}s."